OpenStudy (ammarah):

Find the vertex and find the axis of symmetry for each.. 1. -3x^2 +1 2. y=-4x^ -2x+9

OpenStudy (ammarah):

@Zale101

OpenStudy (anonymous):

@Luigi0210

OpenStudy (ammarah):

plase help.....

OpenStudy (anonymous):

I don't know this

OpenStudy (luigi0210):

Take the derivative

OpenStudy (ammarah):

the wat?

OpenStudy (ammarah):

@aaronq

OpenStudy (ammarah):

@ganeshie8

ganeshie8 (ganeshie8):

easy way is to convert given equation to vertex form :- \(\large a(x-h)^2 + k \) \((h, k )\) is the vertex \(x = h\) is the axis of symmetry

ganeshie8 (ganeshie8):

take first equation :- 1. \(-3x^2 + 1\) is same as, \(-3(x-0)^2 + 1 \)

ganeshie8 (ganeshie8):

simply, compare it wid vertex form, and write down vertex

OpenStudy (ammarah):

wait so the vertext is (0,1)?

ganeshie8 (ganeshie8):

Yup ! good job !!

ganeshie8 (ganeshie8):

once u have vertex, axis of symmetry is simply its x-coordinate

ganeshie8 (ganeshie8):

here, axis of symmetry is the line \(x = 0 \)

OpenStudy (wolf1728):

Taking a derivative is what you learn in calculus. When the derivative equals zero it is a way to find maximums and minimums . 1. -3x^2 +1 2. y=-4x^ -2x+9 Equation 1 -3x^2 +1 The derivative -6x equals zero when x = 0 When x = 0, the 'y' value is y = -3*0 +1 y = 1 So the vertex is at (0, 1) Since the x² coefficient is negative, the graph is concave down and the vertex is a maximum.